OpenStudy (anonymous):

(6/x+1-1/x)/(2/x-4/x+1)

OpenStudy (anonymous):

just to make sure, is it going to be \[\frac{ 6 }{ x+1 }-\frac{ 1 }{ x } / \frac{ 2 }{ x-4 } * \frac{ 1 }{ x+1 }\] ?

OpenStudy (anonymous):

and if i interpreted right, then here's the step

OpenStudy (anonymous):

\[\frac{ 6x-(x+1) }{ (x+1)x } / \frac{ 2 }{ (x-4)(x+1) }\]

OpenStudy (anonymous):

\[\frac{ 5x-1 }{ (x+1)x } * \frac{ (x-4)(x+1) }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ (5x-1)(x-4) }{ 2x }\]